# LLLMs: A Data-Driven Survey of Evolving Research on Limitations of Large Language Models

### 11 Mar 2026 v3

@@ 3.2. Keyword-Based Filtering
This results in a list of 90 keywords (19 unigrams, 44 bigrams, 16 trigrams, and 11 four-grams). Overall, the final keyword set covers key aspects of LLM research (see the full list in Section A of the online supplementary material):
@@ 3.3.1. Human Annotation Task
For papers rated 2–5, annotators highlighted textual evidence pointing to the limitation and, where explicitly expressed in the text, its general type (i.e., the kind of limitation discussed), hereafter referred to as ?evidence?. See Table 2 in the online supplementary material for representative examples of annotated papers and highlighted evidence.
@@ 3.4. Models and Prompting Evaluation
Similar to human annotators highlighting evidence in abstracts to indicate discussions of LLLMs, models are prompted to extract supporting evidence from the text. Each prompt instructs the model: ?Please respond in the following format, providing a rating and supporting evidence for the discussion of LLM limitations in each abstract. Do not include explanations, only cite the evidence found in the abstract.? Prompt 3, included as the most comprehensive, is available in Figure 1 of the online supplementary material.
@@ 3.5. Clustering
To ensure meaningful clusters and minimize spurious outliers, we tune UMAP, HDBSCAN and BERTopic parameters separately for ACL and arXiv. This separation accounts for differences in corpus size and topical heterogeneity. We further apply a distance-based outlier reassignment strategy after observing that many outliers were not true noise but semantically close to existing clusters, suggesting misclassification by HDBSCAN. Full parameter settings and details of the reassignment procedure are provided in Section B of the online supplementary material.
@@ 4.2. LLM and LLLMs Trends Over Time
This suggests a notable shift in NLP research, with LLMs becoming central to the field. In arXiv, growth is more moderate but consistent, hitting just above 30% of papers by the end of the same period. The lower proportion in arXiv might be due to different levels of engagement with LLM research across the categories in our study, as shown in Figure 4 in the online supplementary material. In cs.CL, LLMs are widely discussed, reaching around 80% by early 2025, similar to ACL, while cs.AI shows a sharper but lower rise, peaking around around 50–60%. In contrast, in areas like cs.CV and cs.LG, their presence remains below 20%.
@@
In the following sections, we refine this analysis and examine these emerging discussions in detail through topic clustering. We discuss topics identified with HDBSCAN and LlooM, continue with trends identified with LlooM and provide trends based on HDBSCAN trends in the online supplementary material.
@@ 4.3.1. Topics identified within ACL and arXiv with HDBSCAN and LlooM
Table 7 presents representative papers for topics shared between HDBSCAN and LlooM in the ACL and arXiv datasets. Tables 6 and 7 in the online supplementary material report the dominant terms (top 20 keywords) for each limitation cluster identified with HDBSCAN.
@@
In addition to shared topics, HDBSCAN identifies several clusters that do not emerge under LlooM. Representative papers for these clusters are shown in Table 8 in the online supplementary material. In ACL, the Uncertainty cluster describes behavioral instability, including prompt sensitivity and calibration errors. In arXiv, additional clusters capture a wider range of specialized concerns: Conversational Limitations Code Generation, Healthcare Application, Benchmark Contamination, and Quantization. These clusters reflect both domain-specific challenges (e.g. Healthcare Application, which focuses on the use of LLMs in clinical settings and is primarily concerned with their black-box nature, see example 13) and implementation-level trade-offs. For example, Benchmark Contamination refers to test data leakage into training sets (example 14), and Quantization concerns scaling efficiency (example 15).
@@
A number of topics are specific to LlooM. These are listed below, with explanations based on the prompts used by LlooM to guide topic assignment (see Table 3 and Table 4 in the online supplementary material). For each LlooM-specific cluster, we also provide representative paper examples in Table 8 in the online supplementary material.
  • Trustworthiness (arXiv): the largest cluster in the arXiv set (Figure 7). This category describes concerns about the reliability, transparency, and reproducibility of LLM outputs (see example 19 and 20 which refer to concerns about the reliability of outputs generated by LMs). As a broad category, it often overlaps with related issues like hallucination or alignment, as discussed further in Section D of the online supplementary material.
@@
The comparison of shared and method-specific clusters highlights both similarities and systematic differences between HDBSCAN and LlooM (we quantify these differences in 4.4). While HDBSCAN provides an interpretable, unsupervised clustering of limitations, it assigns each paper to a single cluster and becomes increasingly fragmented as more clusters are added, particularly in the ACL dataset. This makes it difficult to capture overlapping limitations and may complicate the interpretation of trends. By contrast, LlooM allows for broader, non-mutually exclusive categories through multi-label assignment, which reduces fragmentation and provides a more stable basis for longitudinal analysis. Accordingly, we rely on LlooM for the trend analysis in the following section. Trend results derived from HDBSCAN are reported in Section C of the online supplementary material for completeness.
@@ 4.3.2. Trend Analysis
We begin by examining how visible different limitation topics are across the broader LLM research field. Figure 8 shows the annual distribution of LLM limitation topics across ACL and arXiv, normalized by all LLM-focused papers. These proportions reflect the overall visibility of each topic. Additionally, to capture how visibility changes over time, we compute the relative percentage change in LLM-normalized topic share from year $y$ to $y+1$, defined as $\frac{\text{Share}_{y+1}-\text{Share}_{y}}{\text{Share}_{y}}\times 100$, where $\text{Share}_{y}$ refers to the LLM-wide share of a given topic in year $y$, i.e., the proportion of all LLM papers that address that topic (see Table 5 in the online supplementary material).
@@
The stabilization of topic trends coincides with a sharp rise in raw paper counts beginning in 2023 (see Figure 3 in the online supplementary material), indicating not just increased research volume, but a shift toward a more coherent field. Before 2023-Q2, most topics appear in fewer than 25 ACL papers and under 100 in arXiv, making early signals harder to interpret. This growth aligns with the release of ChatGPT in November 2022, as well as the emergence of other major models like GPT-4 (Achiam et al., 2023), PaLM (Chowdhery et al., 2023), and LLaMA (Touvron et al., 2023) between February and July 2023, which likely contributed to the expansion and differentiation of LLM limitations research during this period.
@@ 4.3.3. LLLMs Topics Distribution Across ArXiv Categories
Figure 11. Limitation topic trends across four main arXiv categories (cs.CL, cs.LG, cs.AI, cs.CV). To maintain visual clarity, only the five most frequent topics (based on their overall percentage distribution across categories, as shown in Figure 5 in the online supplementary material) are shown; others are grouped as ?Other.? Grey shading in cs.AI and cs.CV marks periods before 2023 with insufficient data for these categories.
Figure 12. Distribution of limitations-related topics in six arXiv categories with lower paper counts in our dataset, as shown in Figure 5 in the online supplementary material. Each chart includes only papers where the category is assigned as the primary arXiv category. Topics that make up less than 3% are grouped under Other.
Although many categories share topics such as Trustworthiness, Reasoning, Generalization, and Alignment Limitations, both their overall topic composition and temporal dynamics vary by field. Figure 5 (right) in the online supplementary material, shows how topic shares differ across arXiv categories, while Figure 11 illustrates how topics in the four largest categories evolve over time. Key trends in the largest categories are as follows:
@@ 4.4. Clustering Methods Comparison
The aforementioned topic distributions differ depending on the clustering method. In this section, we compare HDBSCAN+BERTopic (Section C in the online supplementary material) and LlooM (4.3.2) to identify which findings are stable and which may be method-specific. We first assess trend agreement, then explore methodological differences to explain any result divergences.
@@ 4.4.1. Cross-Method Comparison Between HDBSCAN and LlooM
To compare the trends, we report (i) Kendall’s Tau for each individual topic (as determined in the trend analyses in Section C in the online supplementary material and 4.3.2) to assess alignment in trend direction and significance, and (ii) the Spearman correlation of the time series between matched topics from HDBSCAN and LlooM to assess the similarity of overall trend shapes. To identify matching topics between HDBSCAN and LlooM, we first identify candidate matches for each cluster based on identical or semantically similar names and validate these by computing the Jaccard overlap between their associated paper sets [^6] All matches except Knowledge Editing were confirmed by highest Jaccard overlap; Knowledge Editing was aligned manually. Specifically, for each cluster produced by HDBSCAN, we compute its Jaccard similarity with all LlooM topics, and vice versa. Given two paper sets $X$ from HDBSCAN and $Y$ from LlooM, the Jaccard similarity is defined as $\text{Jaccard}(X,Y)=\frac{|X\cap Y|}{|X\cup Y|}$. We select the top-1 Jaccard score for each cluster, representing its highest similarity with any topic from the other method. Best-matching topic pairs are shown in Tables 9 and 10 in the online supplementary material, confirming that major topics identified by name also show substantial paper overlap.
@@
These results show that LlooM and HDBSCAN identify similar broad limitation areas but organize papers differently at a finer-grained level. This is supported by Tables 9 and 10 in the online supplementary material, which report the closest topic alignments between methods. Large areas such as Reasoning, Hallucination, Security Risks, and Bias and Fairness appear relatively stable and align well across methods and datasets ($J>0.4$), whereas smaller topics, including Overconfidence and Prompt Sensitivity, are often absorbed into broader categories such as Hallucination. This reflects differences in clustering granularity: LlooM splits topics into overlapping subcategories, while HDBSCAN merges related issues into broader clusters. For example, HDBSCAN merges LlooM’s Security Risks, Privacy Risks, and Trustworthiness into a single Security cluster.
@@ 2. Within LLLMs research, reasoning is the most frequent topic, but research is diverse
Beyond these, LLM limitation research is notably diverse. Our clustering analyses (HDBSCAN and LlooM) reveal a broad spectrum of concerns, ranging from code generation and benchmark contamination to prompt sensitivity and long context. This breadth reflects the current state of LLM limitation research: a fast-growing, methodologically diverse field still defining its major challenges. Additionally, as shown in Section D in the online supplementary material, many papers address multiple limitations simultaneously, reflecting the complexity of emerging concerns.
@@ Appendix A Keyword List
Table 12. Representative examples from the human rating annotation task, with brief explanation for abstracts’ ratings on the 0–5 scale (defined in Table 2 in the main paper). In higher-rated papers (3–5), evidence of LLLMs is highlighted in bold.
@@
Figure 13. Prompt 3 used for LLM-based classification and evidence extraction (as discussed in Section 3.4 of the main paper). Abstract texts and evidence are shortened for brevity.

### 12 Mar 2026 v4

@@ 3.2. Keyword-Based Filtering
This results in a list of 90 keywords (19 unigrams, 44 bigrams, 16 trigrams, and 11 four-grams). Overall, the final keyword set covers key aspects of LLM research (see the full list in Section A supplementary material):
@@ 3.3.1. Human Annotation Task
For papers rated 2–5, annotators highlighted textual evidence pointing to the limitation and, where explicitly expressed in the text, its general type (i.e., the kind of limitation discussed), hereafter referred to as ?evidence?. See Table 12 in the supplementary material for representative examples of annotated papers and highlighted evidence.
@@ 3.4. Models and Prompting Evaluation
Similar to human annotators highlighting evidence in abstracts to indicate discussions of LLLMs, models are prompted to extract supporting evidence from the text. Each prompt instructs the model: ?Please respond in the following format, providing a rating and supporting evidence for the discussion of LLM limitations in each abstract. Do not include explanations, only cite the evidence found in the abstract.? Prompt 3, included as the most comprehensive, is available in Figure 13 of the supplementary material.
@@ 3.5. Clustering
To ensure meaningful clusters and minimize spurious outliers, we tune UMAP, HDBSCAN and BERTopic parameters separately for ACL and arXiv. This separation accounts for differences in corpus size and topical heterogeneity. We further apply a distance-based outlier reassignment strategy after observing that many outliers were not true noise but semantically close to existing clusters, suggesting misclassification by HDBSCAN. Full parameter settings and details of the reassignment procedure are provided in Section B of the supplementary materials.
@@ 4.2. LLM and LLLMs Trends Over Time
This suggests a notable shift in NLP research, with LLMs becoming central to the field. In arXiv, growth is more moderate but consistent, hitting just above 30% of papers by the end of the same period. The lower proportion in arXiv might be due to different levels of engagement with LLM research across the categories in our study, as shown in Figure 16 in the supplementary material. In cs.CL, LLMs are widely discussed, reaching around 80% by early 2025, similar to ACL, while cs.AI shows a sharper but lower rise, peaking around around 50–60%. In contrast, in areas like cs.CV and cs.LG, their presence remains below 20%.
@@
In the following sections, we refine this analysis and examine these emerging discussions in detail through topic clustering. We discuss topics identified with HDBSCAN and LlooM, continue with trends identified with LlooM and provide trends based on HDBSCAN trends in the supplementary material.
@@ 4.3.1. Topics identified within ACL and arXiv with HDBSCAN and LlooM
Table 7 presents representative papers for topics shared between HDBSCAN and LlooM in the ACL and arXiv datasets. Tables 16 and 18 in the supplementary material report the dominant terms (top 20 keywords) for each limitation cluster identified with HDBSCAN.
@@
In addition to shared topics, HDBSCAN identifies several clusters that do not emerge under LlooM. Representative papers for these clusters are shown in Table 19 in the supplementary material. In ACL, the Uncertainty cluster describes behavioral instability, including prompt sensitivity and calibration errors. In arXiv, additional clusters capture a wider range of specialized concerns: Conversational Limitations Code Generation, Healthcare Application, Benchmark Contamination, and Quantization. These clusters reflect both domain-specific challenges (e.g. Healthcare Application, which focuses on the use of LLMs in clinical settings and is primarily concerned with their black-box nature, see example 13) and implementation-level trade-offs. For example, Benchmark Contamination refers to test data leakage into training sets (example 14), and Quantization concerns scaling efficiency (example 15).
@@
A number of topics are specific to LlooM. These are listed below, with explanations based on the prompts used by LlooM to guide topic assignment (see Table 13 and Table 14 in the supplementary material). For each LlooM-specific cluster, we also provide representative paper examples in Table 19 in the supplementary material.
  • Trustworthiness (arXiv): the largest cluster in the arXiv set (Figure 7). This category describes concerns about the reliability, transparency, and reproducibility of LLM outputs (see example 19 and 20 which refer to concerns about the reliability of outputs generated by LMs). As a broad category, it often overlaps with related issues like hallucination or alignment, as discussed further in Section D of the supplementary material.
@@
The comparison of shared and method-specific clusters highlights both similarities and systematic differences between HDBSCAN and LlooM (we quantify these differences in 4.4). While HDBSCAN provides an interpretable, unsupervised clustering of limitations, it assigns each paper to a single cluster and becomes increasingly fragmented as more clusters are added, particularly in the ACL dataset. This makes it difficult to capture overlapping limitations and may complicate the interpretation of trends. By contrast, LlooM allows for broader, non-mutually exclusive categories through multi-label assignment, which reduces fragmentation and provides a more stable basis for longitudinal analysis. Accordingly, we rely on LlooM for the trend analysis in the following section. Trend results derived from HDBSCAN are reported in Section C of the supplementary material for completeness.
@@ 4.3.2. Trend Analysis
We begin by examining how visible different limitation topics are across the broader LLM research field. Figure 8 shows the annual distribution of LLM limitation topics across ACL and arXiv, normalized by all LLM-focused papers. These proportions reflect the overall visibility of each topic. Additionally, to capture how visibility changes over time, we compute the relative percentage change in LLM-normalized topic share from year $y$ to $y+1$, defined as $\frac{\text{Share}_{y+1}-\text{Share}_{y}}{\text{Share}_{y}}\times 100$, where $\text{Share}_{y}$ refers to the LLM-wide share of a given topic in year $y$, i.e., the proportion of all LLM papers that address that topic (see Table 15 in supplementary material).
@@
The stabilization of topic trends coincides with a sharp rise in raw paper counts beginning in 2023 (see Figure 15), indicating not just increased research volume, but a shift toward a more coherent field. Before 2023-Q2, most topics appear in fewer than 25 ACL papers and under 100 in arXiv, making early signals harder to interpret. This growth aligns with the release of ChatGPT in November 2022, as well as the emergence of other major models like GPT-4 (Achiam et al., 2023), PaLM (Chowdhery et al., 2023), and LLaMA (Touvron et al., 2023) between February and July 2023, which likely contributed to the expansion and differentiation of LLM limitations research during this period.
@@ 4.3.3. LLLMs Topics Distribution Across ArXiv Categories
Figure 11. Limitation topic trends across four main arXiv categories (cs.CL, cs.LG, cs.AI, cs.CV). To maintain visual clarity, only the five most frequent topics (based on their overall percentage distribution across categories, as shown in Figure 17 in the supplementary material) are shown; others are grouped as ?Other.? Grey shading in cs.AI and cs.CV marks periods before 2023 with insufficient data for these categories.
Figure 12. Distribution of limitations-related topics in six arXiv categories with lower paper counts in our dataset, as shown in Figure 17. Each chart includes only papers where the category is assigned as the primary arXiv category. Topics that make up less than 3% are grouped under Other.
Although many categories share topics such as Trustworthiness, Reasoning, Generalization, and Alignment Limitations, both their overall topic composition and temporal dynamics vary by field. Figure 17 (right) in the supplementary material, shows how topic shares differ across arXiv categories, while Figure 11 illustrates how topics in the four largest categories evolve over time. Key trends in the largest categories are as follows:
@@ 4.4. Clustering Methods Comparison
The aforementioned topic distributions differ depending on the clustering method. In this section, we compare HDBSCAN+BERTopic (Section C) and LlooM (4.3.2) to identify which findings are stable and which may be method-specific. We first assess trend agreement, then explore methodological differences to explain any result divergences.
@@ 4.4.1. Cross-Method Comparison Between HDBSCAN and LlooM
To compare the trends, we report (i) Kendall’s Tau for each individual topic (as determined in the trend analyses in Section C in the supplementary material and 4.3.2) to assess alignment in trend direction and significance, and (ii) the Spearman correlation of the time series between matched topics from HDBSCAN and LlooM to assess the similarity of overall trend shapes. To identify matching topics between HDBSCAN and LlooM, we first identify candidate matches for each cluster based on identical or semantically similar names and validate these by computing the Jaccard overlap between their associated paper sets [^6] All matches except Knowledge Editing were confirmed by highest Jaccard overlap; Knowledge Editing was aligned manually. Specifically, for each cluster produced by HDBSCAN, we compute its Jaccard similarity with all LlooM topics, and vice versa. Given two paper sets $X$ from HDBSCAN and $Y$ from LlooM, the Jaccard similarity is defined as $\text{Jaccard}(X,Y)=\frac{|X\cap Y|}{|X\cup Y|}$. We select the top-1 Jaccard score for each cluster, representing its highest similarity with any topic from the other method. Best-matching topic pairs are shown in Table 20 and Table 21 in the supplementary material, confirming that major topics identified by name also show substantial paper overlap.
@@
These results show that LlooM and HDBSCAN identify similar broad limitation areas but organize papers differently at a finer-grained level. This is supported by Tables 20 and 21 in the supplementary material, which report the closest topic alignments between methods. Large areas such as Reasoning, Hallucination, Security Risks, and Bias and Fairness appear relatively stable and align well across methods and datasets ($J>0.4$), whereas smaller topics, including Overconfidence and Prompt Sensitivity, are often absorbed into broader categories such as Hallucination. This reflects differences in clustering granularity: LlooM splits topics into overlapping subcategories, while HDBSCAN merges related issues into broader clusters. For example, HDBSCAN merges LlooM’s Security Risks, Privacy Risks, and Trustworthiness into a single Security cluster.
@@ 2. Within LLLMs research, reasoning is the most frequent topic, but research is diverse
Beyond these, LLM limitation research is notably diverse. Our clustering analyses (HDBSCAN and LlooM) reveal a broad spectrum of concerns, ranging from code generation and benchmark contamination to prompt sensitivity and long context. This breadth reflects the current state of LLM limitation research: a fast-growing, methodologically diverse field still defining its major challenges. Additionally, as shown in Section D in the supplementary materials, many papers address multiple limitations simultaneously, reflecting the complexity of emerging concerns.
@@ Appendix A Keyword List
Table 12. Representative examples from the human rating annotation task, with brief explanation for abstracts’ ratings on the 0–5 scale (defined in Table 2). In higher-rated papers (3–5), evidence of LLLMs is highlighted in bold.
@@
Figure 13. Prompt 3 used for LLM-based classification and evidence extraction (as discussed in 4.3.3). Abstract texts and evidence are shortened for brevity.